• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name bogdanw

    Log entry time 00:11:30 on February 1,2009

    Entry number 258514

    keyword=shift summary

    smooth production
    18:23 Stop for target test/tuning
    19:17 Resume Production after target laser problem is fixed.
    The shift is very smooth but target polarization drop to 60%